OpenStudy (anonymous):

I need help with the intermediates steps of a sum of squares problem. I have a trinomial and the answers it factors into but cannot work it backwards correctly

OpenStudy (anonymous):

a sum of squares will have complex numbers in the factorized form

OpenStudy (anonymous):

the squar is always equal for example: For the first one: The goal of these problems is to get the left hand side of the equation to be a perfect square and then take the square root of both sides and solve for answers. x^2+16x+64=81 The left side of the equation factors into (x+8)^2 so taking the square root of both sides x+8=+ or - 9 so x= 1 or x= -17

OpenStudy (anonymous):

are you talking about factoring a sum of squares or completing the square?

OpenStudy (anonymous):

can i get a medal plzzzzzzzzzzzzzzzzzzz

OpenStudy (anonymous):

if y^2-2xy-x^2=0 and the answers are y=(-1+,- 2^.5)x what are the intermediate steps working backward from those answers?

OpenStudy (anonymous):

you have to try to get the left side of the equation to a perfect square then you square root both sides then solve

OpenStudy (anonymous):

medal plzzz

OpenStudy (anonymous):

sum of squares! \[f(1)=1^2+f(0)\] \[f(2)=2^2+f(1)\] . . . \[f(n)=n^2+(f(n-1))\] Sum all the expressions above: \[\sum_{1}^{n}(x^n)=f(n)-f(0)\] \[f(n)=ax^3+bx^2+cx+d\] \[\sum_{1}^{n}(x^n)=ax^3+bx^2+cx+d\] this is the notion you need. then you compare polynomials and get to the answer. i think it is \[f(n)=(n^3)/6+(n^2)/2+n/3\]

OpenStudy (anonymous):

just dunno if I get the question right! hehe

OpenStudy (anonymous):

thanks both of you but I need to work backward from the answer

OpenStudy (anonymous):

to rearrive at y^2-2xy-x^2=0?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I think I am totally off topic haha, can't wait to see the answer

OpenStudy (anonymous):

\[y^2-2xy-x^2=0\] \[(y-x)^2 -2x^2 = 0\] \[y - x = \pm (\sqrt{2})x\] \[y = x \pm (\sqrt{2})x = (1 \pm \sqrt{2})x\]

OpenStudy (anonymous):

thanks but I'm timed out on my library computer

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

y^2−2xy−x^2=0 (y−x)^2 − 2x^2=0 y−x=±(2^0.5)x y=x±(2^0.5)x=(1±(2^0.5))x

OpenStudy (anonymous):

i think some people got confused when you called it "sum of squares" instead of "completing the square" :)

OpenStudy (anonymous):

this works as y^2 - 2xy -x^2 = y^2 - 2xy + x^2 -2x^2 = (y-x)^2 - 2x^2